OpenStudy (anonymous):

Am I correct? Question: Explain what the discriminant is and what it tells you about the solutions of a quadratic equation.

OpenStudy (anonymous):

A discriminant is the expression inside the square root. If discriminant is positive then the square root is real there are two real solutions but when the discriminant is zero, the square root is also zero and there is one unique real solution. If the equation has two roots that are identical in value, then it has a 'double root' at that value. Lastly, if the discriminant is negative then the square root is imaginary, and the original equation has two distinct complex roots.

OpenStudy (anonymous):

Am I correct on this?

OpenStudy (anonymous):

Yes..:D You are correct.

OpenStudy (unklerhaukus):

for \(ax^2+bx+c=0\) The discriminant is \[\Delta=b^2-4ac\] yeah, you might want to say the the square-root you are referring to is the the square root in the quadratic formula, \(x=\frac{-b\pm\sqrt{\Delta}}{2a}\) what you have said it true you may also want to inclue if the discriminant is a square number; the roots will be rational